OpenStudy (anonymous):

Verify the identity. sin 4x = 2 sin 2x cos 2x

OpenStudy (anonymous):

Start with sin(4x)=sin(2x+2x)

OpenStudy (anonymous):

yeah.........whats next??

OpenStudy (anonymous):

@campbell_st i don't understand..

OpenStudy (anonymous):

Let \(y=2x\) \[\begin{split} \sin (4x) &= \sin (2y) &\text{sub } y=2x\\ &= \sin (y+y) \\ &= \sin (y)\cos(y)+\sin(y) \cos (y)\\ &=2 \sin (y)\cos(y)&\text{sub }y=2x\\ &= 2 \sin (2x) \cos (2x) \end{split}\]

OpenStudy (campbell_st):

ok... so there is a trig identity called the sum of 2 angles for sin its \[\sin(a + b) = \sin(a)\cos(b) + \cos(a)(\sin(b)\] you will need to use it. so in your question split the 4x in 2 equal parts 2x and 2x so \[\sin(4x) = \sin(2x + 2x)\] now using the expansion above you will get \[\sin(2x + 2x) = \sin(2x)\times \cos(2x) + \cos(2x) \times \sin(2x)\] it will simplify to \[2\sin(2x)\cos(2x)\] I hope this helps.

OpenStudy (anonymous):

@campbell_st it helps so much thank you!!

OpenStudy (campbell_st):

glad it makes sense. good luck